OpenStudy (anonymous):

find the equation of the curve which passes through the pts (2,-2) and (4,2) and for which dy/dx =x^2(x-k) where k is a constant.

OpenStudy (anonymous):

is there a graph with this question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i actually don't know this but i will see if any of my friends know this and help

OpenStudy (amistre64):

integrate, and use the intial values to define the curve

OpenStudy (amistre64):

dy/dx =x^2(x-k) dy =x^2(x-k) dx \[\int~dy =\int~x^2(x-k) dx\]

OpenStudy (anonymous):

i did i cant the right answer.

OpenStudy (anonymous):

what about the k?

OpenStudy (amistre64):

what was your integration result?

OpenStudy (amistre64):

the k is some constant, you are given 2 points to initialize so that it forms a system of 2 equations in 2 unknowns

OpenStudy (amistre64):

let k=3 for the moment if its distracting you

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

when i tried to integrate with the k ,i got 1/4(x^4) -1/3 k (x^3)

OpenStudy (amistre64):

\[\int~dy =\int~x^2(x-k) dx\] \[y =\int~x^3-kx^2~dx\] \[y =\frac14x^4-\frac13kx^3+C\] initialize it with the points (2,-2) and (4,2) \[-2 =\frac142^4-\frac k32^3+C\] \[2 =\frac144^4-\frac k34^3+C\] solve for k and C

OpenStudy (anonymous):

ok i get it know

OpenStudy (amistre64):

by elimination, -1 the top and add \[2 =-\frac142^4+\frac k32^3-C\]\[2 =\frac144^4-\frac k34^3+C\] \[4=\frac14(4^4-2^4)+k(\frac13(2^3-4^3))\] \[4-\frac14(4^4-2^4)=k(\frac13(2^3-4^3))\] \[\frac{4-\frac14(4^4-2^4)}{\frac13(2^3-4^3)}=k\]

OpenStudy (anonymous):

k=3 and then i replace in the other eqn to find C....

OpenStudy (anonymous):

tysm i got the answer right :)

OpenStudy (amistre64):

:) youre welcome